### v2

# Two Hornich-Hlawka-type and Gram matrix-based inequalities

Abstract
This paper deals with Hornich-Hlawka-type and Gram matrix-based inequalities in inner product spaces (IPS). On the one hand, the Hornich-Hlawka inequality, or Hlawka’s inequality, or quadrilateral inequality, is an inequality concerning three vectors in a quadrilateral. Namely, it states that for all $x,y,z$, $\lVert x\rVert+\lVert y\rVert+\lVert z\rVert+\lVert x+y+z\rVert\geq\lVert x+y\rVert+\lVert x+z\rVert+\lVert y+z\rVert$. On the other hand, inequalities based on Gram matrices are very important because they generate generically all inequalities implying three vectors in an IPS. In the first part of our paper, we introduce a stronger version of Hornich-Hlawka inequality, namely, for all vectors $x,y,z$, $\|x\|\|y\|+\|z\|\|x+y+z\|\geq\|x+z\|\|y+z\|$. In the second part of our paper, we establish a sharp inequality based on Gram matrices that implies, in particular, the standard Cauchy-Schwarz inequality.
00footnotetext: 2020 Mathematics Subject Classification. 46C05; 26D15 00footnotetext: Key words and phrases. Hornich-Hlawka inequality; Gram matrix; Cauchy-Schwarz inequality; inner product spaces; norm inequalities
1 Introduction
On the one hand, the celebrated Hornich-Hlawka inequality [2, 3, 1, 4, 5, 6, 7, 8, 9, 10, 11] in normed vector spaces $E$ states
$\forall x,y,z\in E:\lVert x\rVert+\lVert y\rVert+\lVert z\rVert+\lVert x+y+z\rVert\geq\lVert x+y\rVert+\lVert x+z\rVert+\lVert y+z\rVert$
While it has been proved that this inequality fails to hold in some normed vector spaces, it does hold, nonetheless, in a large class of normed vector spaces including inner product spaces.
On the other hand, important inequalities can be established from the nonnegativity of Gram matrices. Actually, all inequalities implying three vectors in an IPS can be generically derived from the nonnegativity of the Gram matrices of these vectors.
After this introduction, our article is divided into 2 other sections. In section 2, we prove a stronger form of Hornich-Hlawka inequality in inner product spaces, namely: for all vectors $x,y,z$, $\|x\|\|y\|+\|z\|\|x+y+z\|\geq\|x+z\|\|y+z\|$. In section 3, we prove a sharp inequality based on Gram matrices that strengthens, in particular, the Cauchy-Schwarz inequality. Hereafter, $H$ denotes a real inner product space, with scalar product $\langle\cdot,\cdot\rangle$ and norm $\lVert\cdot\rVert$.
2 A stronger Hornich-Hlawka inequality
Theorem 2.1 (Strong Hornich–Hlawka inequality).
For all $x,y,z\in H$, we have
$\lVert x\rVert\lVert y\rVert+\lVert z\rVert\lVert x+y+z\rVert\geq\lVert x+z\rVert\lVert y+z\rVert,$
Moreover, equality holds if and only if $x,y,z$ are linearly dependent and their dependence coefficients satisfy one of the equivalent conditions listed in proposition 2.3
Remark 2.2.
  • When $H=\mathbb{R}$ or $\mathbb{C}$, the inequality can be easily proved by expanding both sides and using the multiplicativity of the modulus (this property is senseless when we consider vector norms instead of the modulus of complex numbers).
  • This stronger Hornich-Hlawka inequality implies the classical Hornich-Hlawka inequality. Indeed, since both sides of the classical Hornich-Hlawka inequality are nonnegative, we may replace them with their squares.
    For the left-hand side, we obtain
    $\bigl(\|x\|+\|y\|+\|z\|+\|x+y+z\|\bigr)^{2}=2\bigl(\|x\|^{2}+\|y\|^{2}+\|z\|^{2}\bigr)+2\bigl(\langle x,y\rangle+\langle y,z\rangle+\langle z,x\rangle\bigr)$
    $2\bigl(\|x\|+\|y\|+\|z\|\bigr)\|x+y+z\|+2\bigl(\lVert x\rVert\lVert y\rVert+\lVert x\rVert\lVert z\rVert+\lVert y\rVert\lVert z\rVert\bigr).$
    For the right-hand side, we have
    $\bigl(\|x+y\|+\|y+z\|+\|z+x\|\bigr)^{2}=2\bigl(\|x\|^{2}+\|y\|^{2}+\|z\|^{2}\bigr)+2\bigl(\langle x,y\rangle+\langle y,z\rangle+\langle z,x\rangle\bigr)$
    $2\bigl(\|x+y\|\cdot\|y+z\|+\|y+z\|\cdot\|z+x\|+\|z+x\|\cdot\|x+y\|\bigr).$
    Thus, the classical Hornich-Hlawka inequalityis is equivalent to
    $\bigl(\|x\|+\|y\|+\|z\|\bigr)\|x+y+z\|+\bigl(\lVert x\rVert\lVert y\rVert+\lVert x\rVert\lVert z\rVert+\lVert y\rVert\lVert z\rVert\bigr).$
    $\geq\|x+y\|\cdot\|y+z\|+\|y+z\|\cdot\|z+x\|+\|z+x\|\cdot\|x+y\|$
    which, after regrouping, is equivalent to
    $\bigl(\lVert x\rVert\lVert y\rVert+\|z\|\cdot\|x+y+z\|-\|y+z\|\cdot\|z+x\|\bigr)+$
    $\bigl(\lVert y\rVert\lVert z\rVert+\|x\|\cdot\|x+y+z\|-\|z+x\|\cdot\|x+y\|\bigr)+$
    $\bigl(\lVert x\rVert\lVert z+\|y\|\cdot\|x+y+z\|-\|x+y\|\cdot\|y+z\|\bigr)\geq 0.$
    Therefore, it is clear that the stronger Hornich-Hlawka inequality implies the classical Hornich-Hlawka inequality.
Proof.
Idea: we reduce the inequality to a quartic form in the Gram matrix entries and show that its minimum over the PSD cone is attained on the boundary.
  • First step. Reduction.
    Denote by $L$ and $R$ the left and right hand sides of this inequality, respectively.
    We have
    $L^{2}$ $=\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\left(\lVert x\rVert^{2}+\lVert y\rVert^{2}+\lVert z\rVert^{2}+2\langle x,y\rangle+2\langle x,z\rangle+2\langle y,z\rangle\right)+2\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert$
    $=\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\lVert x\rVert^{2}+\lVert z\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{4}+2\lVert z\rVert^{2}\langle x,y\rangle+2\lVert z\rVert^{2}\langle x,z\rangle+2\lVert z\rVert^{2}\langle y,z\rangle$
    $\quad\quad+2\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert$
    and
    $R^{2}$ $=\left(\lVert x\rVert^{2}+\lVert z\rVert^{2}+2\langle x,z\rangle\right)\left(\lVert y\rVert^{2}+\lVert z\rVert^{2}+2\langle y,z\rangle\right)$
    $=\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\lVert x\rVert^{2}+\lVert z\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{4}+2\langle x,z\rangle\lVert y\rVert^{2}+2\langle x,z\rangle\lVert z\rVert^{2}$
    $\quad\quad+2\langle y,z\rangle\lVert x\rVert^{2}+2\langle y,z\rangle\lVert z\rVert^{2}+4\langle x,z\rangle\langle y,z\rangle$
    We have
    $L\geq R$
    $\Leftrightarrow L^{2}\geq R^{2}$
    $\Leftrightarrow\frac{L^{2}-\left(\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\lVert x\rVert^{2}+\lVert z\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{4}\right)}{2}\geq\frac{R^{2}-\left(\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\lVert x\rVert^{2}+\lVert z\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{4}\right)}{2}$
    $\Leftrightarrow\lVert z\rVert^{2}\langle x,y\rangle+\lVert z\rVert^{2}\langle x,z\rangle+\lVert z\rVert^{2}\langle y,z\rangle+\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert\geq$
    $\quad\quad\quad\langle x,z\rangle\lVert y\rVert^{2}+\langle x,z\rangle\lVert z\rVert^{2}+\langle y,z\rangle\lVert x\rVert^{2}+\langle y,z\rangle\lVert z\rVert^{2}+2\langle x,z\rangle\langle y,z\rangle$
    $\Leftrightarrow\lVert z\rVert^{2}\langle x,y\rangle+\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert\geq\lVert x\rVert^{2}\langle y,z\rangle+\lVert y\rVert^{2}\langle x,z\rangle+2\langle x,z\rangle\langle y,z\rangle$
    This inequality is equivalent to
    $\Leftrightarrow\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert\geq\lVert x\rVert^{2}\langle y,z\rangle+\lVert y\rVert^{2}\langle x,z\rangle-\lVert z\rVert^{2}\langle x,y\rangle+2\langle x,z\rangle\langle y,z\rangle$
    Denote by L and R the left and right hand sides of this last inequality, respectively.
    If $\textbf{R}\leq 0$, the inequality is trivial, so we can suppose that $\textbf{R}\geq 0$.
    Set
    $a=\lVert x\rVert,\quad b=\lVert y\rVert,\quad c=\lVert z\rVert,\quad p=\langle x,y\rangle,\quad q=\langle x,z\rangle,\quad r=\langle y,z\rangle.$
    We have
    $\textbf{L}^{2}$ $=\lVert x\rVert^{2}\lVert y\rVert^{2}\lVert z\rVert^{2}\left(\lVert x\rVert^{2}+\lVert y\rVert^{2}+\lVert z\rVert^{2}+2\langle x,y\rangle+2\langle x,z\rangle+2\langle y,z\rangle\right)$
    $=a^{2}b^{2}c^{2}(a^{2}+b^{2}+c^{2}+2p+2q+2r).$
    and
    $\textbf{R}^{2}$ $=\lVert x\rVert^{4}\langle y,z\rangle^{2}+\lVert y\rVert^{4}\langle x,z\rangle^{2}+\lVert z\rVert^{4}\langle x,y\rangle^{2}+4\langle x,z\rangle^{2}\langle y,z\rangle^{2}$
    $\quad+2\lVert x\rVert^{2}\lVert y\rVert^{2}\langle x,z\rangle\langle y,z\rangle-2\lVert x\rVert^{2}\lVert z\rVert^{2}\langle x,y\rangle\langle y,z\rangle-2\lVert y\rVert^{2}\lVert z\rVert^{2}\langle x,y\rangle\langle x,z\rangle$
    $\quad+4\lVert x\rVert^{2}\langle x,z\rangle\langle y,z\rangle^{2}+4\lVert y\rVert^{2}\langle y,z\rangle\langle x,z\rangle^{2}-4\lVert z\rVert^{2}\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle$
    $=a^{4}r^{2}+b^{4}q^{2}+c^{4}p^{2}+4q^{2}r^{2}+2a^{2}b^{2}qr-2a^{2}c^{2}pr-2b^{2}c^{2}pq+4a^{2}qr^{2}+4b^{2}rq^{2}-4c^{2}pqr$
    Consider
    $\xi$ $=\textbf{L}^{2}-\textbf{R}^{2}$
    $=a^{2}b^{2}c^{2}(a^{2}+b^{2}+c^{2}+2p+2q+2r)$
    $\quad-\left(a^{4}r^{2}+b^{4}q^{2}+c^{4}p^{2}+4q^{2}r^{2}+2a^{2}b^{2}qr-2a^{2}c^{2}pr-2b^{2}c^{2}pq+4a^{2}qr^{2}+4b^{2}rq^{2}-4c^{2}pqr\right)$
  • Second step. The minimum of $\xi$ over the PSD cone is attained on the boundary.
    Consider $\xi$ as a quadratic function in one variable, say $p$, for fixed $a,b,c,q,r$ (satisfying the relevant conditions, see next).
    The coefficient of $p^{2}$ is $c^{4}\leq 0$, so it is a downward-opening parabola.
    The set of allowed values $a,b,c\geq 0$, $p,q,r$ are such that the Gram matrix
    $G=\begin{pmatrix}a^{2}&p&q\\ p&b^{2}&r\\ q&r&c^{2}\end{pmatrix}$
    is positive semidefinite.
    This requires all principal minors to be nonnegative: $|p|\leq ab$, $|q|\leq ac$, $|r|\leq bc$, and $\det G\geq 0$.
    Since the allowed range for $p$ is the interval where $\det G\geq 0$ and $|p|\leq ab$, which is closed and bounded, the minimum of $\xi$ over this interval occurs at one of the endpoints.
    The endpoints correspond to boundaries where either $\det G=0$ or $p=\pm ab$ (but note that $p=\pm ab$ implies the 2x2 minor for $x,y$ is zero, which implies $x,y$ are linearly dependent, which forces $\det G=0$). Thus, the minima occur on the boundary where $G$ is singular ($\det G=0$).
    To show $\xi\geq 0$ on this boundary, assume linear dependence:
    • If $z=\lambda x+\mu y$, then substitute
      $c^{2}$ $=\lambda^{2}a^{2}+\mu^{2}b^{2}+2\lambda\mu p$
      $q$ $=\lambda a^{2}+\mu p$
      $r$ $=\lambda p+\mu b^{2}$
      After substitution and simplification, $\xi$ factors as
      $\xi=(a^{2}b^{2}-p^{2})\left(a^{2}\lambda(\lambda+1)-b^{2}\mu(\mu+1)\right)^{2}.$
      The first factor is nonnegative by the Cauchy-Schwarz inequality, and the second factor is a square, hence nonnegative. Therefore, $\xi\geq 0$.
    • If $x=\lambda y+\mu z$, then substitute
      $a^{2}$ $=\lambda^{2}b^{2}+\mu^{2}c^{2}+2\lambda\mu r,$
      $p$ $=\lambda b^{2}+\mu r,$
      $q$ $=\lambda r+\mu c^{2}.$
      After substitution and simplification, $\xi$ factors as
      $\xi=\Bigl(b^{2}c^{2}-r^{2}\Bigr)\Bigl(b^{2}\lambda(\lambda+1)+c^{2}\mu(\mu+1)+2\lambda(\mu+1)r\Bigr)^{2}.$
      The first factor is nonnegative by the Cauchy-Schwarz inequality, and the second factor is a square, hence nonnegative. Therefore, $\xi\geq 0$.
    • If $y=\lambda x+\mu z$, then substitute
      $b^{2}$ $=\lambda^{2}a^{2}+\mu^{2}c^{2}+2\lambda\mu q,$
      $p$ $=\lambda a^{2}+\mu q,$
      $r$ $=\lambda q+\mu c^{2}.$
      After substitution and simplification, $\xi$ factors as
      $\xi=\left(a^{2}c^{2}-q^{2}\right)\left(a^{2}\lambda(\lambda+1)+c^{2}\mu(\mu+1)+2\lambda(\mu+1)q\right)^{2}.$
      The first factor is nonnegative by the Cauchy-Schwarz inequality, and the second factor is a square, hence nonnegative. Therefore, $\xi\geq 0$.
    Therefore, $\xi\geq 0$ on the boundary.
  • Conclusion.
    Since $\xi\geq 0$ on the boundary, we have $\xi\geq 0$ throughout the domain.
    Equality arises in configurations where the vectors are linearly dependent, with additional conditions on the scalar products and norms: the inequality becomes an equality when the geometry is "flat" (vectors in a plane or a line).
Proposition 2.3 (Equality cases).
Let $x,y,z\in H$. Equality holds in the inequality
$\|x\|\|y\|+\|z\|\|x+y+z\|=\|x+z\|\|y+z\|$
if and only if the vectors $x,y,z$ are linearly dependent and one of the following cases occurs.
  1. (i)
    There exist real numbers $\lambda,\mu$ such that
    $z=\lambda x+\mu y$
    and
    $\|x\|^{2}\,\lambda(\lambda+1)=\|y\|^{2}\,\mu(\mu+1).$
  2. (ii)
    There exist real numbers $\lambda,\mu$ such that
    $x=\lambda y+\mu z$
    and
    $\|y\|^{2}\,\lambda(\lambda+1)+\|z\|^{2}\,\mu(\mu+1)+2\lambda(\mu+1)\langle y,z\rangle=0.$
  3. (iii)
    There exist real numbers $\lambda,\mu$ such that
    $y=\lambda x+\mu z$
    and
    $\|x\|^{2}\,\lambda(\lambda+1)+\|z\|^{2}\,\mu(\mu+1)+2\lambda(\mu+1)\langle x,z\rangle=0.$
In all cases, the Gram matrix
$G=\begin{pmatrix}\|x\|^{2}&\langle x,y\rangle&\langle x,z\rangle\\ \langle x,y\rangle&\|y\|^{2}&\langle y,z\rangle\\ \langle x,z\rangle&\langle y,z\rangle&\|z\|^{2}\end{pmatrix}$
is singular. In particular, equality occurs only in flat configurations, i.e., when the vectors $x,y,z$ lie in a common one or two dimensional subspace.
3 An inequality based on Gram matrices
Theorem 3.1.
We have, for all vectors $x,y,z\in H$ and real numbers $\alpha,\beta,\gamma$:
$\alpha^{2}\|x\|^{2}\langle y,z\rangle^{2}+\beta^{2}\|y\|^{2}\langle x,z\rangle^{2}+\gamma^{2}\|z\|^{2}\langle x,y\rangle^{2}+2(\alpha\beta+\alpha\gamma+\beta\gamma)\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle\geq 0.$
Proof.
Given $x,y,z$ in $H$, the Gram matrix associated to $(x,y,z)$ is
$G=\begin{pmatrix}\langle x,x\rangle&\langle x,y\rangle&\langle x,z\rangle\\ \langle y,x\rangle&\langle y,y\rangle&\langle y,z\rangle\\ \langle z,x\rangle&\langle z,y\rangle&\langle z,z\rangle\end{pmatrix}$.
It is positive semidefinite, so for any real numbers $\alpha,\beta,\gamma$, we have
$Q(\alpha,\beta,\gamma)=(\alpha,\beta,\gamma)\,G\,\begin{pmatrix}\alpha\\ \beta\\ \gamma\end{pmatrix}\geq 0.$
Expanding, we have
$Q(\alpha,\beta,\gamma)=\alpha^{2}\|x\|^{2}+\beta^{2}\|y\|^{2}+\gamma^{2}\|z\|^{2}+2\alpha\beta\langle x,y\rangle+2\alpha\gamma\langle x,z\rangle+2\beta\gamma\langle y,z\rangle\geq 0.$
Replace
$\alpha\mapsto\alpha\langle y,z\rangle,$
$\beta\mapsto\beta\langle x,z\rangle,$
$\gamma\mapsto\gamma\langle x,y\rangle.$
Then
$R(\alpha,\beta,\gamma)$ $=Q(\alpha\langle y,z\rangle,\beta\langle x,z\rangle,\gamma\langle x,y\rangle)$
$=\alpha^{2}\|x\|^{2}\langle y,z\rangle^{2}+\beta^{2}\|y\|^{2}\langle x,z\rangle^{2}+\gamma^{2}\|z\|^{2}\langle x,y\rangle^{2}$
$\quad+2(\alpha\beta+\alpha\gamma+\beta\gamma)\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle\geq 0.$
Corollary 3.2.
  • If $\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle>0$, we have
    $\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+\|z\|^{2}\langle x,y\rangle^{2}\geq 3\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle.$
  • If $\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle<0$, we have
    $\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+\|z\|^{2}\langle x,y\rangle^{2}\geq-6\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle.$
These inequalities are sharp.
Remark 3.3.
  • This corollary can be understood as a reverse inequality of the Gram determinant inequality
    $\det G\geq 0$
    which translates to
    $\|x\|^{2}\|y\|^{2}\|z\|^{2}+2\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle\geq\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+\|z\|^{2}\langle x,y\rangle^{2}.$
  • The first point of the corollary implies the Cauchy-Schwarz inequality, and therefore all other inequalities deduced from it. One can show this by taking $z=y$.
Proof.
  • If
    $\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle\geq 0$
    Then R(1,1,2)>0R(1,1,-2)>0 implies
    $\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+4\|z\|^{2}\langle x,y\rangle^{2}\geq 6\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle.$
    Summing up symmetrically and dividing by 6, we get
    $\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+\|z\|^{2}\langle x,y\rangle^{2}\geq 3\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle.$
  • If
    $\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle<0$
    Then $R(1,1,1)\geq 0$ implies
    $\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+\|z\|^{2}\langle x,y\rangle^{2}\geq-6\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle.$
Sharpness. For $D>0$, take $x=y=z$ nonzero. It yields LHS = RHS.
For $D<0$, take $x=(1,0,0)$, $y=\left(\frac{1}{2},\frac{\sqrt{3}}{2},0\right)$, $z=\left(\frac{1}{2},-\frac{\sqrt{3}}{2},0\right)$.
We have
$\langle x,y\rangle=1\cdot\frac{1}{2}+0\cdot\frac{\sqrt{3}}{2}=\frac{1}{2}$
$\langle x,z\rangle=1\cdot\frac{1}{2}+0\cdot\left(-\frac{\sqrt{3}}{2}\right)=\frac{1}{2}$
$\langle y,z\rangle=\frac{1}{2}\cdot\frac{1}{2}+\frac{\sqrt{3}}{2}\cdot\left(-\frac{\sqrt{3}}{2}\right)=\frac{1}{4}-\frac{3}{4}=-\frac{1}{2}$
So
$RHS=-6\left(\frac{1}{2}\right)\left(\frac{1}{2}\right)\left(-\frac{1}{2}\right)=\frac{6}{8}=\frac{3}{4}.$
Moreover
$\|x\|^{2}\langle y,z\rangle^{2}=1\cdot\frac{1}{4}=\frac{1}{4}$
$\|y\|^{2}\langle x,z\rangle^{2}=1\cdot\frac{1}{4}=\frac{1}{4}$
$\|z\|^{2}\langle x,y\rangle^{2}=1\cdot\frac{1}{4}=\frac{1}{4}$
So
$LHS=\frac{1}{4}+\frac{1}{4}+\frac{1}{4}=\frac{3}{4}.$
Equality holds. ∎